# Is Compression Really Linear with Code Intelligence?

### v4

@@ 4.1 Setup
Multiple state-of-the-art Code LLMs are adopted, covering model scales from 1.5B to 70B parameters. Selected models are listed as follows: Qwen2.5-Coder (hui2024qwen2, ), DeepSeek-Coder (guo2024deepseek, ), StarCoder (li2023starcoder, ; lozhkov2024starcoder, ), CodeGemma (team2024codegemma, ), CodeLlama (roziere2023code, ), Llama3.1 (grattafiori2024llama, ), Yi-Coder (yicoder, ), Granite-Code (mishra2024granite, ), and OpenCoder (huang2024opencoder, ).
@@
Our format annealing training set consists of 4,995 QA pairs. We use DeepSpeed ZeRO-3 optimization with BF16 precision on 8*A100-80GB GPU. The batch size, learning rate, and warm-up ratio are set to 128, 1e-4, and 0.1, respectively. Each models are trained for up to 4 epochs. After each epoch, the ratio of successful extraction from outputs is evaluated. When it exceeds 99%, the training is stopped and the current checkpoint is used for the OmniCodeBench evaluation.
@@
$N$ denotes the number of code task categories, $M_{k}$ represents the benchmarks in task $k$, $w_{k,j}=\frac{1}{N}\cdot\frac{|D_{k,j}|}{\sum_{i=1}^{M_{k}}|D_{k,i}|}$ is the hierarchical weight for benchmark $j$ in task $k$, and $s_{k,j}\in[0,1]$ is the performance on benchmark $j$.This entropy-based weighting strategy ensures initial task fairness via the $\frac{1}{N}$ term and a data-driven benchmark emphasis via the $|D_{k,j}|$ proportion. After extensive experiments comparing the fitting curves of ${C}$ and $\log{C}$, We choose $\log{C}$ as our final metric, which represents code intelligence. Relevant experiments are presented in Appendix.
4.2 Results
@@
We consider that this logarithmic relationship aligns more in line with theoretical expectations. From the perspective of model capabilities, when benchmark difficulty is appropriately moderate, models with high BPC values lack sufficient knowledge to solve problems effectively, whereas those with relatively low BPC exhibit substantial evaluation metric improvements even with marginal BPC reductions. This nonlinear correspondence strongly rationalize the observed logarithmic law.
@@
5 Discussion
Figure 5: Fitting curves of four different code tasks (code generation, fix, reasoning and explanation). In each figure, benchmarks unrelated to the current task are discarded. The values of code intelligence are pre-processed in logarithmic scale.
@@ 6 Conclusion
In this paper, we explore the logarithmic relationship between code BPC and code intelligence, presenting a more refined approach to evaluating the capabilities of code-focused LLMs. We introduce a sliding window evaluation technique to ensure fair comparisons across models with varying context window lengths, addressing potential biases in BPC calculation. To further strengthen our evaluation, we develop OmniCodeBench, a multilingual, multi-task benchmark, alongside a Format Annealing training methodology. This approach improves pattern adherence while preserving the intrinsic knowledge of foundational models, effectively minimizing format-related errors that commonly affect model outputs. Our findings provide new possibilities for investigating the relationship between compression and intelligence of LLMs.
Limitation
However, our focus remains primarily on coder-oriented LLMs, with limited exploration of general natural language LLMs. Additionally, the language distribution in OmniCodeBench remains a heavy focus on Python. Future work should address these limitations and extend the established BPC-intelligence correlation beyond code-related domains.

### v5

@@ 4.1 Setup
Multiple state-of-the-art Code LLMs are adopted, covering model scales from 1.5B to 70B parameters. Selected models are listed as follows: Qwen2.5-Coder [26], DeepSeek-Coder [21], StarCoder [34, 37], CodeGemma [52], CodeLlama [50], Llama3.1 [19], Yi-Coder [1], Granite-Code [40], and OpenCoder [23].
@@
Our format annealing training set consists of 4,995 QA pairs. We use DeepSpeed ZeRO-3 optimization with BF16 precision on 8*A100-80GB GPU. The batch size, learning rate [31], and warm-up ratio are set to 128, 1e-4, and 0.1, respectively. Each models are trained for up to 4 epochs. After each epoch, the ratio of successful extraction from outputs is evaluated. When it exceeds 99%, the training is stopped and the current checkpoint is used for the OmniCodeBench evaluation.
@@
NN denotes the number of code task categories, $M_{k}$ represents the benchmarks in task $k$, $w_{k,j}=\frac{1}{N}\cdot\frac{|D_{k,j}|}{\sum_{i=1}^{M_{k}}|D_{k,i}|}$ is the hierarchical weight for benchmark $j$ in task $k$, and $s_{k,j}\in[0,1]$ is the performance on benchmark $j$.This entropy-based weighting strategy ensures initial task fairness via the $\frac{1}{N}$ term and a data-driven benchmark emphasis via the $|D_{k,j}|$ proportion. After extensive experiments comparing the fitting curves of C{C} and $\log{C}$, We choose $\log{C}$ as our final metric, which represents code intelligence. Relevant experiments are presented in the supplementary material.
4.2 Results and Findings
@@
We think that this logarithmic relationship aligns more in line with theoretical expectations. From the perspective of model capabilities, when benchm ark difficulty is appropriately moderate, models with high BPC lack sufficient knowledge to solve problems effectively, whereas those with relatively low BPC exhibit substantial evaluation metric improvements even with marginal BPC reductions. This nonlinear correspondence strongly rationalizes the observed logarithmic law.
@@
5 Discussions
Figure 5: Fitting curves of four different code tasks (code generation, fix, reasoning and explanation). In each figure, benchmarks unrelated to the current task are discarded. The values of code intelligence are in logarithmic scale.
@@ 5 Discussions
Abnormal impact from benchmarks: During a single benchmark fitting process, we observe that McEval struggles to establish logarithmic relationships. Further analysis of the outputs reveal that on one hand, models with abnormally low McEval scores consistently produce lengthy invalid continuations containing various repetitions, meaningless characters, or multilingual noise. This issue appears most severe in Qwen-series models, causing their scores to drop to single-digit levels. On the other hand, models with unusually high McEval scores are disproportionately boosted by their C/C++ performance. For instance, OpenCoder-1.5B significantly outperforms OpenCoder-8B solely due to its superior C/C++ metrics, while both models exhibit normal Python performance. The inherent difficulty and linguistic diversity of McEval make it challenging to serve as a standalone reference point for observing our fitting logarithmic relationship.
@@
Insight for further research of LLM: Based on and inspired by our conclusions, this approach enables more precise modeling of the relationship between model compression and intelligence in code-related domains and even across general fields. This advancement not only facilitates rapid performance estimation for smaller models but also allows predictive scaling for larger-parameter models, thereby avoiding substantial redundant computational expenditure.
@@ 6 Conclusion
This paper explores the logarithmic relationship between code BPC and code intelligence. To this end, we introduce a sliding window evaluation technique to ensure fair comparisons across models with varying context window lengths, addressing potential biases in BPC calculation. To further strengthen our evaluation, we develop OmniCodeBench, a multilingual, multi-task benchmark, alongside a Format Annealing training methodology. This approach improves pattern adherence while preserving the intrinsic knowledge of foundational models, effectively minimizing format-related errors that commonly affect model outputs. Our findings provide new possibilities for investigating the relationship between compression and intelligence of LLMs.
Appendix A Data Synthesis Pipeline for Training Set
The training set integrates three primary sources: CommitPackFT [41], CodeSearchNet [27], and LeetCode, each subjected to a specialized preprocessing workflow. The CommitPackFT data containing paired “old contents” and “new contents” code revisions serves as the foundation for the code fix task. From an initial pool of approximately 50,000 sampled instances, we implemented a multistage cleaning protocol: first removing entries where old and new contents are identical or contain empty inputs, followed by filtering out files with non-executable placeholders (e.g., “TODO”, “FIXME”). We subsequently leveraged GPT-4o to automatically verify that the code in old content reliably triggers run-time errors while their corresponding new content resolve these issues, preserving associated error logs for prompt generation. A final manual inspection phase eliminates ambiguous or meaningless cases, resulting in 3,855 rigorously validated samples. These curated instances are further standardized using GPT-4o to match the input-output formatting specifications of LiveCodeBench code fix benchmark, ensuring seamless integration with downstream evaluation frameworks.
The LeetCode dataset comprises 3,077 programming challenges and their corresponding solutions scraped from LeetCode international platform, exhibiting inherently high data quality due to its curated nature. To clean the data, we implemented a standardized filtering protocol: (1) excluding video-based solutions; (2) randomly selecting a solution for each problem; (3) structurally aligning each instance with LiveCodeBench’s code generation task format with the assistance of GPT-4o. The standardized prompt template integrates four key components: the problem statement, illustrative examples (including input and output cases), algorithmic constraints and the initial function signature, thereby preserving the essential context required for reproducible code generation tasks.
The CodeSearchNet dataset extracts function docstrings as instructions, with a high degree of structural repetition. We removed entries where the output character length exceeds 5,000, as well as those with too short or meaningless instructions. The data was then sorted in descending order by length, and 5-gram deduplication was applied to removing any instructions or output containing “TODO” or “FIXME”. We used GPT-4o to filter suitable data for each language, and finally, a comprehensive sampling was performed to obtain 8,699 data samples. Ultimately, we sampled 4,995 training data samples from three data sources as our training set.
Appendix B Why Choose LogC
Figure 7: Linear relationship between BPC and code intelligence on our open-source validation set.
Figure 7 shows that with a RMSE of 0.115 the linear fitting curve performs worse than the logarithmic relationship curve with a RMSE of only 0.057. Therefore, in our paper the values of code intelligence are all pre-processed in logarithmic scale.
Appendix C Individual Fitting Curve of Each Benchmark
The fitting results of each benchmark are displayed in Figure 8,  LABEL:fig:2,  LABEL:fig:3,  9,  10,  11,  12,  13,  14,  15 which seperately represent HumanEval, BigCodeBench-Full, BIRD, Spider, LiveCodeBench-Fix, CRUXEval-I, CRUXEval-O and HumanEvalExplain.
Figure 8: Fitting curve between HumanEval benchmark and its corresponding metric.
Figure 9: Fitting curve between BigCodeBench-Full benchmark and its corresponding metric.
Figure 10: Fitting curve between BIRD benchmark and its corresponding metric.
Figure 11: Fitting curve between Spider benchmark and its corresponding metric.
Figure 12: Fitting curve between LiveCodeBench-Fix benchmark and its corresponding metric.
Figure 13: Fitting curve between CRUXEval-I benchmark and its corresponding metric.
Figure 14: Fitting curve between CRUXEval-O benchmark and its corresponding metric.
Figure 15: Fitting curve between HumanEvalExplain benchmark and its corresponding metric.
Appendix D Details of Open-Source Validation Set
Detailed data quantity and proportion of our open-source validation set are shown in Table 2.
Table 2: More details of our open-source validation set.
Language Number of Data Entries Number of Characters Proportion of Characters
Python 1, 500 7, 086, 432 41.21%
Go 500 1, 598, 829 9.30%
R 300 1, 460, 408 8.49%
C 100 714, 577 4.16%
C++ 200 861, 125 5.01%
Java 500 2, 020, 641 11.75%
JavaScript 500 1, 866, 524 10.85%
HTML 100 892, 826 5.19%
SQL 200 695, 673 4.05%
Total 3, 900 17, 197, 035 100.00%